OpenStudy (anonymous):

Rose worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer.

OpenStudy (anonymous):

This is the problem.

OpenStudy (texaschic101):

um....I don't think I will be able to help because the problem doesn't look wrong to me......so obviously I am doing something wrong as well ???? :( @satellite73 ......I feel stupid asking for some help on this easy problem....

OpenStudy (anonymous):

looks correct

OpenStudy (texaschic101):

Is it the multiplying of the two, because it is still in brackets, and you would do that before dividing....wouldn't you ?

OpenStudy (texaschic101):

@skullpatrol ....can you take a peek at this....

OpenStudy (anonymous):

its like asking what she did wrong but I don't see what she did wrong

OpenStudy (texaschic101):

neither do I

OpenStudy (texaschic101):

@genius12 ....can you look at this problem....its confusing me

OpenStudy (texaschic101):

@mathslover ....I feel stupid for not knowing this....can you look at it ?

OpenStudy (texaschic101):

ok...step by step......30/ 5 (3 - 1)- 4 do parenthesis first... 30/5(2) - 4 still parenthesis 30/10 - 4 division 3 - 4 -1........what am I doing wrong ?

OpenStudy (texaschic101):

oh...wait.....30/5 * 2 - 4 division first 6 * 2 - 4 12 - 4 6 Is that it ?

OpenStudy (anonymous):

yup. that's it

OpenStudy (texaschic101):

oh...I feel stupid for not catching that the first couple of times

OpenStudy (anonymous):

\[30/5(3-1) -4 = \frac{ 30 }{ 5 } (3 -1) - 4\]

OpenStudy (texaschic101):

did you get that @Tayloresc

mathslover (mathslover):

First divide, 30/5 = 6 and then multiply with 2. 12- 4 = 8

mathslover (mathslover):

Line 3 has the mistake

OpenStudy (anonymous):

Thanks everyone!!!!

mathslover (mathslover):

@texaschic101 , 12-4 = 8 not 6:)

OpenStudy (texaschic101):

oops....my bad